OpenStudy (anonymous):

Show sin 4θ = 4sin θ cos θ cos 2 θ

OpenStudy (anikay):

do you know double angle formulas?

OpenStudy (anonymous):

not really because m confused with it

OpenStudy (anikay):

but you have been 'taught' them in class?

OpenStudy (anonymous):

im teaching myself right now

OpenStudy (anikay):

http://www.sosmath.com/trig/Trig5/trig5/trig5.html this can help

OpenStudy (anonymous):

okay, but what do i do to with the whole equation

OpenStudy (anonymous):

im really confused with this part

OpenStudy (anikay):

so how I would do this is to split up the sin(4theta) into this \[\sin(2x) \] where x= 2theta

OpenStudy (anikay):

if you want me to keep going just say so, but after every step I do I'm going to give you the chance to pick it up and do it yourself.

OpenStudy (anonymous):

so what will the new equation look like if we use sin instead of cos

OpenStudy (anikay):

I'm manipulating only the left side. that's how these kinds of problems work, you only mess with one side.

OpenStudy (anikay):

I'll type out a step or two in order.

OpenStudy (anikay):

\[\sin(4\theta)\] which we then alter to become \[\sin(2x)\] then we use the double angle formula to get 2sin(x)cos(x) then we remember that x = 2theta and rewrite it like this \[2\sin(2\theta)\cos(2\theta)\] got it from here?

OpenStudy (anonymous):

would we equal it to 1 and bring it to the other side

OpenStudy (anonymous):

would sin 4θ=4sinθcosθcos 2θ become sin2 θ=2sin(2θ)cos(2θ)

OpenStudy (anonymous):